A car move at a constant velocity of 20m/s for 100seconds and then accelerated uniformly to a velocity of 30m/s over 15 seconds. Calculate the distance covered during the motion


1
Expert's answer
2021-06-11T11:21:18-0400

The distance covered at the first part of the motion:


"d_1 = 20m/s\cdot 100s = 2000m"

By definition, the acceleration at the second part of the motion:


"a_2 = \dfrac{30m/s - 20m/s}{15s} = \dfrac23m/s^2"


The distance covered at the second part of the motion:


"d_2 = 20m/s\cdot 15 s + \dfrac{2/3m/s^2\cdot (15s)^2}{2} = 375m"

The total distance is:


"d = d_1 + d_2 = 2375m"

Answer. 2375m.